OpenStudy (anonymous):

1over3(24p-66)=3p+43

OpenStudy (amistre64):

Lover3?

OpenStudy (anonymous):

Its not a mixed fraction

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

legend at 100, does it kick at 101? or is that just it?

OpenStudy (anonymous):

\[\frac{1}{3}(24p-66)=8p-22\]

OpenStudy (anonymous):

dunno

OpenStudy (anonymous):

yes that equation is right

OpenStudy (anonymous):

solve using that

OpenStudy (anonymous):

i think it said on more level to go but this one doesn't move

OpenStudy (amistre64):

congratz on the lvl100 regardless tho ;)

OpenStudy (anonymous):

can you answer guys

OpenStudy (anonymous):

thanks. actually this might be it for me for a while. busy starting wedensday

OpenStudy (amistre64):

patience max

OpenStudy (anonymous):

can you answer the question first please

OpenStudy (amistre64):

multiply it all by 3 to clear the fraction

OpenStudy (amistre64):

the it becomes simple addition to work out til you divide of whats left on the p

OpenStudy (anonymous):

\[8p-22=3p+43\] first a) subtract 3p from both sides b) add 22 to both sides c) divide both sides by 5

OpenStudy (amistre64):

busy starting wednesday sounds like a good thing :) i was not busy at all yesterday and went about stircrazy ...

OpenStudy (anonymous):

didn't see you here yesterday at all

OpenStudy (amistre64):

no internet and library was closed ..

OpenStudy (anonymous):

oooooh. how is school coming along?

OpenStudy (amistre64):

its creeping along at a snails pace. grades are fine, just at my age the material feels like kindergarten all over again :)

OpenStudy (anonymous):

lol i bet time for coffee have fun oh and i trust you got the book ok yes?

OpenStudy (amistre64):

yep, got it; and thank you again ;)

OpenStudy (anonymous):

@satellite - i managed to get a secondhand edition of Spivak's Calculus

OpenStudy (amistre64):

\[\frac{1}{3}(24p-66)=8p-22\] when we multiply it all by 3 we can clear the fraction .. \[(3)\frac{1}{3}(24p-66)=3(8p-22)\]\[24p-66=24p-66\] from here it looks lke p can be any value it wants to be